OpenStudy (anonymous):

plz help i need answer as soon as possible :) A boat traveled 12 miles with the current and 15 miles against it. The boat's average speed is 10 miles per hour. A. Let c represent the current. Write a rational expression for the time with the current. B. Let c represent the current. Write a rational expression for the time against the current. C. Write an equation to represent the total time. D. Simplify the equation written in Part C. E. Find the total time if the speed of the current

OpenStudy (anonymous):

Let x be the speed of boat in still waters and c be the speed of current. DOWNSTREAM ------------------------- Speed = x + c mph A) Time taken = 12 / ( x+c) hours B) Time taken = 15/ ( x-c) hours C) Total time = 12/ (x+c) + 15/ ( x-c) hours Average speed = (( x+c) + ( x-c) ) / 2 = 2x /2 = x = 10 mph D) 12/ (10+ c) + 15/ ( 10-c) 12 ( 10-c) + 15( 10+c) ----------------------------------- 100 - c^2 = (270 +3c )/ ( 100- c^2) E) 279/ 91 = 3.066 hours

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

Babe Here You Go general formula : v = s / t or t = s / v but the current modify the real speed, so going with the current v is really v+c, and against is v-c A. t = s / v = 12 / (10 + c) B. t = s / v = 15 / (10 - c) C Total time = 12 / (10+c) + 15 / (10 - c) D. Tt = 12 / (10+c) + 15 / (10 - c) = (12*(10 - c) + 15 * (10 + c)) / (10 + c)*(10 - c) Tt = (120 - 12c + 150 + 15c) / (10^2 - c^2) Tt = (270 + 3c ) / (100 - c^2) E Tt = (270 + 3c ) / (100 - c^2) = (270 + 3* 3 ) / (100 - 3^2) = (270 + 9 ) / (100 - 9) = 279 / 91 Tt = 3.066 h

OpenStudy (anonymous):

no thank you lol lol